OpenStudy (anonymous):

Find all solutions of the form infinite sum from n=0 to infinity AnX^(n), find the differential equation to x^(2)y''-6xy'+10y=0

OpenStudy (anonymous):



OpenStudy (nikvist):

\[x^2y''-6xy'+10y=0\]\[y=\sum\limits_{n=0}^{+\infty}a_n\cdot x^n\quad,\quad y'=\sum\limits_{n=1}^{+\infty}a_n\cdot nx^{n-1}\quad,\quad y''=\sum\limits_{n=2}^{+\infty}a_n\cdot n(n-1)x^{n-2}\]\[x^2\sum\limits_{n=2}^{+\infty}a_n\cdot n(n-1)x^{n-2}-6x\sum\limits_{n=1}^{+\infty}a_n\cdot nx^{n-1}+10\sum\limits_{n=0}^{+\infty}a_n\cdot x^n=0\]\[\sum\limits_{n=2}^{+\infty}a_n\cdot n(n-1)x^{n}+\sum\limits_{n=1}^{+\infty}-6a_n\cdot nx^{n}+\sum\limits_{n=0}^{+\infty}10a_n\cdot x^n=0\]\[10a_0+4a_1x+\sum\limits_{n=2}^{+\infty}a_n(n-5)(n-2)x^n=0\]\[\Rightarrow\quad y=a_5x^5+a_2x^2\]